OpenStudy (anonymous):

Write the equation of the quadratic function with roots -2 and 4 and a vertex at (1, 9).

OpenStudy (welshfella):

the vertex form of a quadratic is a(x - h)^2 + k where the vertex is at (h,k)

OpenStudy (welshfella):

so in your case we have f(x) = a(x - 1)^2 + 9

OpenStudy (welshfella):

now you need to find the value of a which you can do using the given roots

OpenStudy (anonymous):

so what is a?

OpenStudy (welshfella):

a is a number to be found

OpenStudy (welshfella):

the roots are -2 and 4 so points of intersection on x-axis are (-2,0) and (4,0) So plugging these values into the above equation and solving for a will give you the final equation

OpenStudy (anonymous):

I'm really not good at this. but i think i'm kinda getting it? could you just show me please?

OpenStudy (welshfella):

ok f(x) = 0 when a - -2 so 0 = a(-2-1)^2 + 9 0 = 9a + 9 -9a = 9 a = -1 so we;ve found a

OpenStudy (anonymous):

so the answer is f(x) = -1(x - 1)^2 + 9?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thanks so much. i will fan and medal you. :)

OpenStudy (welshfella):

they might want it in standard form though just expand and simplify

OpenStudy (welshfella):

yw ty

OpenStudy (welshfella):

they haven't asked for that so leaving in in the above form should be ok

OpenStudy (anonymous):

yep, that should be fine. thanks again.